If r,s,t are prime numbers and p,q are the positive integers such that L.c.m of p,q is r2s4t2., then the total number of ordered pairs(p,q) is ??

Grade:11

2 Answers

Abhishek Singh
9 years ago
no. of pais possible=3*5*3=45
 
Abhishek Singh
9 years ago
sorry it is incorrect solution, i mistakened it for factors rather than lcm question
Correct solution
Selection for p &q
 p                                      q                     ways
r0                                                 r2                                   1
r1                                                  r2                                  1
r2                                                   r0 ,r1,   r2                      1
 
Total way to select r=5
s0                             s4                                            1 way
s1                             s4                                            1 way 
s2                             s4                                            1 way
s3                              s4                                           1 way
s4                             s0,s1,s3,s4                     5 way
 taota ways to select r=9
similarly for t=5
total ways = 5*9*5= 225